OpenStudy (ilovecake):

Which equation is a related equation for this equation? 18/a =7 HELP ME PLEASE A. a= 18 times 7 B. a= 18/7 C. a+7=18 D. a= 7/18

OpenStudy (ilovecake):

MEDAL

OpenStudy (ilovecake):

So I do 18 divided by 7 which is 2

OpenStudy (ilovecake):

Is that right?

OpenStudy (anonymous):

That's not right. \[\frac{ 18 }{ a }=7\] First you need to clear the fraction. So multiply both sides by a. That gives you: 18 = 7a Can you solve that for a? One more step.

OpenStudy (ilovecake):

I think the answer would be D?

OpenStudy (anonymous):

No. If you have 18 = 7a, you divide by the number in front of the variable. So you'd divide both sides by 7...

OpenStudy (adamk):

You had it. 18/7. So which letter says "a = 18/7"?

OpenStudy (ilovecake):

B

OpenStudy (ilovecake):

I mean Thank for explaining it. Wow sorry. I made a mistake by saying that.